Recent Log Entries

RF Health Check

Log Number: 4130423Logbook(s): Tags: Body: 

Beam Conditions:
Total Linac Current (R2XXITOT) = 230.91
Hall A Current (IBC1H04CRCUR2) = 9.69
Hall B Current (IPM2C24A.IENG) = 4.37913
Hall C Current (IBC3H00CRCUR4) = 38.43
Hall D Current (IBCAD00CRCUR6) = 309
BSY Current (IBC4C00CRCUR5) = 0
INJ Energy (MMSINJEGAIN) = 118.148
NL Energy (MMSLIN1EGAIN) = 1047
SL Energy (MMSLIN2EGAIN) = 1047

Gradient Problems:
1L03-2 GLDER=373 (>=200) with RF OFF - Check Detuning
1L06-1 Gradient Loop Detector Error RMS (GLDER)=81 (<0 or >=80ppm)
1L07-1 Local Oscillator Power (LOPW)=14.9507 (<=70mW)
1L07-2 Local Oscillator Power (LOPW)=16.6881 (<=70mW)
1L07-4

iCalibrate: Hall A

Log Number: 4130412Logbook(s): Body: Calibration ParametersHall:ATarget: Carbon HolePass: 4Note: Calibrated On:2023-Jan-14 23:45Calibrated By: anthonyFile:A-2023-01-14_2345_Carbon_Hole_Pass-4.hcdIon ChamberExisting Setpoing (rads/hr)New Setpoint (rads/hr)Compton1,998<

iCalibrate: Hall D

Log Number: 4130395Logbook(s): Body: Calibration ParametersHall:DTarget: JD70-103 50um 45/135 degPass: 5.5Note: Calibrated On:2023-Jan-22 13:58Calibrated By: schoeneFile:D-2023-01-22_1358_JD70-103_50um_45_135_deg_Pass-5.5.hcdIon ChamberExisting Setpoing (rads/hr)New Setpoint (rads/hr)

Lowered Chopper 2 Temp Setpoint

Log Number: 4130394Logbook(s): Entry Makers: michaelmBody: 

Lowered Chopper 2's temperature setpoint on the Watlow controler 0.2 Degrees C from 43.9 to 43.7. Did this in response to the Buncher high temperature alarm. Other parameters (Grad, FWD PWR, REF PWR, Phase) have been steady. The temperature change raised Chopper 2's X&Y REF PWR but the Buncher REF PWR dropped.

Master Oscillator

Log Number: 4130384Logbook(s): Entry Makers: mdiazBody: 

We measured the LO (1427 MHz) and the IF (70 MHz) on a bypassed cavity in 2L26 for spectrum content and phase noise. All within specification.

1427 MHz: PN < 100 mdeg, No obvious spectral content out to 5 MHz

70 MHz: PN < 5 mdeg, No obvious spectral content out to 5 MHz except for the MOMOD signal at 390 Hz and that was -83 dBc which is within specification.

Pages